Show that X^2+7x-44=0


1
Expert's answer
2021-11-10T18:35:09-0500

x2+7x-44=0

Rewriting the expression by writing 7x as the difference. It will be like:

x2+11x-4x-44=0

Factor the expression by factoring out x from the expression

x*(x+11)-4(x+11)=0

Separate into possible cases

(x+11)*(x-4)=0

Solve the equations:

x+11=0

x-4=0

The equation has 2 solutions:

1) x=-11

2) x=4

Checking:

1) (-11)2+7*(-11)-44=121-77-44=0

2) 42+7*4-44=16+28-44=0